### v2

# On endomorphism algebras of silting complexes over hereditary abelian categories

Abstract.
Let $\mathcal{E}$ be the class of finite-dimensional algebras isomorphic to endomorphism algebras of silting complexes over hereditary abelian categories. It is proved that the class $\mathcal{E}$ is closed under taking idempotent quotients, idempotent subalgebras and $\tau$-reduction. We also show that the proper class consisting of shod algebras is also closed under these operations. In addition, several classic classes of algebras—including laura, glued, weakly shod algebras—are proved to be closed under idempotent quotients, thereby generalizing a known result originally established for specific idempotents.
Key words and phrases:
silting reduction, $\tau$-reduction, quasi-silted algebra, idempentent quotient
2020 Mathematics Subject Classification:
16G10, 16E10, 18E30
This work was supported by the National Natural Science Foundation of China (Grant No. 12571040)
1. Introduction
For a given algebra $A$, the study of its idempotent subalgebra $eAe$ and the idempotent quotient $A/AeA$ associated with an idempotent $e\in A$ is a central theme in representation theory (see, for instance, [CPS, AC, Xi, Xu, CK]). A primary motivation for focusing on these two constructions originates from the theory of recollements of derived categories. It is well known that under suitable conditions, the bounded derived category $\mathcal{D}^{b}(A)$ can be “glued” from $\mathcal{D}^{b}(eAe)$ and $\mathcal{D}^{b}(A/AeA)$ via a recollement [CPS]. This framework provides a powerful mechanism for homological reduction, allowing one to decompose the homological properties of AA—such as the finiteness of its global dimension—into those of $eAe$ and $A/AeA$ [AKLY].
From this perspective, establishing the closure properties of a specific class of algebras under these operations is of fundamental importance. It transforms the class into a self-contained system where global conjectures and structural problems can be tackled through recursive arguments. For example, the fact that the class of $2$-Calabi-Yau tilted algebras arising from hereditary categories is closed under idempotent quotients was a key step in proving the connectedness of cluster-tilting graphs for hereditary categories [FG]. Similarly, the proof of the connectedness of $\tau$-tilting graphs for gentle algebras is based on the fact that the class of gentle algebras is closed under $\tau$-reduction [FGLZ].
In the representation theory of finite-dimensional algebras, the study of endomorphism algebras of canonical objects—such as tilting modules, support $\tau$-tilting modules, and (2-term) silting complexes—is a central theme. These constructions define several fundamental classes of algebras, including tilted, quasi-tilted, silted, and quasi-silted algebras. A landmark result by Happel, Reiten, and Smalø [HRS] established that almost hereditary algebras are precisely the endomorphism algebras of tilting objects in hereditary abelian categories. This categorical characterization was further extended by Buan and Zhou [BZa, BZb], who demonstrated that shod algebras can be realized as endomorphism algebras of 2-term silting complexes over hereditary abelian categories. As a further generalization, endomorphism algebras of silting complexes over a finite-dimensional hereditary algebras were investigated in [ALLT] from the view point of structure of categories.
The purpose of this work is to investigate the class $\mathcal{E}$ of finite-dimensional algebras realized as endomorphism algebras of basic silting complexes over hereditary abelian categories, specifically focusing on their behavior under the operations of taking idempotent subalgebras and idempotent quotients. Employing the machinery of silting reduction in triangulated categories, we develop an approach to analyze the stability of this class of algebras. Our main results, which are interconnected and proved using this method, establish the following closure properties.
Theorem 1.1 (Theorem 3.3, Theorem 3.5).
Let $A\in\mathcal{E}$ and $e\in A$ be an idempotent element. Then
  • (1)
    $A/AeA\in\mathcal{E}$. Moreover, if $A$ is quasi-silted, then so is $A/AeA$.
  • (2)
    $eAe\in\mathcal{E}$. Moreover, if $A$ is quasi-silted (resp. quasi-tilted), then so is $eAe$.
We remark that the final assertion of (2) was previously established in [AC]. However, our proof is entirely different from the one provided there.
Note that the idempotent quotient can be viewed as a special case of the $\tau$-reduction introduced by Jasso [J]. It is therefore natural to consider whether the class $\mathcal{E}$ is closed under the more general operation of $\tau$-reduction.
Theorem 1.2 (Theorem 3.9).
Let $A\in\mathcal{E}$ and $Z$ be a $\tau$-rigid $A$-module.
  • (1)
    The $\tau$-reduction of $A$ with respect to $Z$ belongs to $\mathcal{E}$.
  • (2)
    If $A$ is quasi-silted, then the $\tau$-reduction of $A$ with respect to $Z$ is also quasi-silted.
The proof of Theorem 1.1, which establishes the closure property under idempotent quotients, naturally motivates a further investigation into the classical classes of algebras studied by Assem and Coelho [AC]. Specifically, we are led to consider whether these well-known classes—namely laura, glued, weakly shod, and shod algebras—are also closed under quotients by arbitrary idempotents. Utilizing an independent analysis with techniques distinct from silting reduction, we provide an affirmative answer to this question.
Theorem 1.3 (Theorem 4.1).
Let $A$ be a finite-dimensional $k$-algebra and $e\in A$ be an idempotent element.
  1. (1)
    If $A$ is a laura algebra, then so is $A/AeA$.
  2. (2)
    If $A$ is a right (or left) glued algebra, then so is $A/AeA$.
  3. (3)
    If $A$ is a weakly shod algebra, then so is $A/AeA$.
  4. (4)
    If $A$ is a shod algebra, then so is $A/AeA$.
This result significantly generalizes the work of Zito [Z], who proved a similar statement for specific idempotents.
The paper is organized as follows. We provide the necessary preliminaries and the main method of silting reduction in Section 2. In Section 3, we present the proof of Theorems 1.1-1.2. Section 4 is devoted to the proof of Theorem 1.3. In Section 5, we provide some examples clarifying the boundaries of our results.
Notation
Throughout this paper, let $k$ be a field. By a hereditary abelian category, we mean a $\operatorname{\mathsf{Hom}}$-finite and $\operatorname{\mathsf{Ext}}$-finite $k$-linear hereditary abelian categry.
  • For a category $\mathcal{C}$ and a subcategory $\mathcal{X}$ of $\mathcal{C}$. We denote by $\mathcal{X}^{\perp}$ the full subcategory of $\mathcal{C}$ consisting of objects $X\in\mathcal{C}$ such that $\operatorname{\mathsf{Hom}}_{\mathcal{C}}(\mathcal{X},X)=0$. We define $^{\perp}\mathcal{X}$ dually.
  • We denote by $[1]$ the suspension functor of a triangulated category $\mathcal{T}$. For subcategories $\mathcal{X}$ and $\mathcal{Y}$ of $\mathcal{T}$, we denote by $\mathcal{X}*\mathcal{Y}$ the full subcategory consisting of objects $E\in\mathcal{T}$ such that there is a triangle $X\to E\to Y\to X[1]$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. Given an object $X$ we denote by $\operatorname{\mathsf{thick}}X$ the smallest thick subcategory of $\mathcal{T}$ containing $X$.
  • In a Krull-Schmidt category, for an object $X$, denote by $|X|$ the number of pairwise non-isomorphic indecomposable direct summands of $X$, and by $\operatorname{\mathsf{add}}\nolimits X$ the full subcategory of the finite direct sums of direct summands of $X$.
  • For an algebra $A$, we denote by $\operatorname{\mathsf{mod}}\nolimits A$ its category of finitely generated right modules, and by $\operatorname{\mathsf{ind}}A$ the set of isomorphism classes of indecomposable right $A$-modules. For an element $x\in A$, write $\langle x\rangle=AeA$.
  • For a module $M\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by $\operatorname{\mathsf{Fac}}M$ the full subcategory of factor modules of direct sum of finite copies of $M$. Dually, denote by $\operatorname{\mathsf{Sub}}M$ the full subcategory of the submodules of direct sum of finite copies of $M$.
2. Preliminaries
2.1. Perpendicular subcategories in triangulated categories
Let $\mathcal{T}$ be a triangulated category and $\mathcal{X}$ a subcategory of $\mathcal{T}$. A morphism $f:X\to Y$ is called a right $\mathcal{X}$-approximation of $Y\in\mathcal{T}$, if $X\in\mathcal{X}$ and $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},f):\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},X)\to\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},Y)$ is surjective. The subcategory $\mathcal{X}$ is contravariantly finite in $\mathcal{T}$, if each object in $\mathcal{T}$ has a right $\mathcal{X}$-approximation. Dually, we can define left $\mathcal{X}$-approximation and covariantly finite subcategories. The subcategory $\mathcal{X}$ is called functorially finite, if it is both contravariantly finite and covariantly finite.
Lemma 2.1.
[IO, Proposition 5.33] Let $\mathcal{X}$ and $\mathcal{Y}$ be full subcategories of $\mathcal{T}$ .
  • (1)
    If $\mathcal{X}$ and $\mathcal{Y}$ are contravariantly finite in $\mathcal{T}$ , then so is $\mathcal{X}\ast\mathcal{Y}$.
  • (2)
    If $\mathcal{X}$ and $\mathcal{Y}$ are covariantly finite in $\mathcal{T}$ , then so is $\mathcal{X}*\mathcal{Y}$.
Recall that a pair of subcategories $(\mathcal{X},\mathcal{Y})$ of $\mathcal{T}$ is a torsion pair of $\mathcal{T}$, if $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},\mathcal{Y})=0$ and $\mathcal{X}*\mathcal{Y}=\mathcal{T}$. The following useful result is known as Wakamatsu’s Lemma (see [AI, Lemma 2.22]).
Lemma 2.2.
Let $\mathcal{M}$ be a subcategory of $\mathcal{T}$ such that $\mathcal{M}*\mathcal{M}\subseteq\mathcal{M}$.
  • (1)
    If $\mathcal{M}$ is contravariantly finite, then $(\mathcal{M},\mathcal{M}^{\perp})$ is a torsion pair of $\mathcal{T}$;
  • (2)
    If $\mathcal{M}$ is covariantly finite, then $(\!^{\perp}\mathcal{M},\mathcal{M})$ is a torsion pair of $\mathcal{T}$.
Let $\mathcal{S}$ be a thick subcategory of $\mathcal{T}$, denote by $\mathcal{T}/\mathcal{S}$ the Verdier quotient of $\mathcal{T}$ with respect to $\mathcal{S}$. Denote by $\mathbb{L}:\mathcal{T}\to\mathcal{T}/\mathcal{S}$ the localization functor. The following is also well-known, see [DF, Lemma 2.2] for instance.
Lemma 2.3.
Let $\mathcal{S}$ be a thick subcategory of $\mathcal{T}$.
  1. (1)
    If $\mathcal{S}$ is contravariantly finite, then the restriction $\mathcal{S}^{\perp}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence,
  2. (2)
    If $\mathcal{S}$ is covariantly finite, then the restriction $^{\perp}\mathcal{S}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence,
2.2. Silting theory
Let $\mathcal{T}$ be a $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulated category. An object $T\in\mathcal{T}$ is called a presilting object if $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(T,T[>0])=0$. A presilting object $T\in\mathcal{T}$ is silting if $\operatorname{\mathsf{thick}}T=\mathcal{T}$. The following result gives a characterization of a thick subcategory generated by a presilting object.
Lemma 2.4.
[AI, Proposition 2.17] Let $D\in\mathcal{T}$ be a presilting object. Then
$\operatorname{\mathsf{thick}}D=\bigcup\limits_{l\geq 0}\operatorname{\mathsf{add}}\nolimits(D[-l])*\operatorname{\mathsf{add}}\nolimits(D[-l+1])*\cdots*\operatorname{\mathsf{add}}\nolimits(D[l]).$
There are triangulated categories which do not have silting objects. But if $\mathcal{T}$ admits a silting object, then the Grothendieck group $K_{0}(\mathcal{T})$ of $\mathcal{T}$ is a free abelian group of rank $|T|$ ([AI, Theorem 2.27]). In particular, each silting object of $\mathcal{T}$ has the same number of pairwise non-isomorphic indecomposable direct summands.
Let $T=M\oplus\overline{T}$ be a basic silting object of $\mathcal{T}$. Consider the triangle
$M\xrightarrow{f_{M}}T_{M}\to N\to M[1],$
where $f_{M}$ is a minimal left $\operatorname{\mathsf{add}}\nolimits\overline{T}$-approximation of $M$. According to [AI, Theorem 2.31], $N\oplus\overline{T}$ is a basic silting object of $\mathcal{T}$ and $\mu_{M}^{-}(T):=N\oplus\overline{T}$ is called the left mutation of $T$ with respect to $M$.
We denote by $\operatorname{\mathsf{silt}}\mathcal{T}$ the set of isomorphism classes of basic silting objects of $\mathcal{T}$.
Theorem 2.5.
[AI, Theorem 2.37] Let $\mathcal{T}$ be a Krull-Schmidt triangulated category, $\mathcal{S}$ a functorially finite thick subcategory of $\mathcal{T}$ and $\mathcal{T}/\mathcal{S}$ the Verdier quotient. Denote by $\mathbb{L}:\mathcal{T}\to\mathcal{T}/\mathcal{S}$ the localization functor. For any $D\in\operatorname{\mathsf{silt}}\ \mathcal{S}$, there is a bijective map
$\{T\in\operatorname{\mathsf{silt}}\ \mathcal{T}~|~D\in\operatorname{\mathsf{add}}\nolimits T\}\to\operatorname{\mathsf{silt}}\ \mathcal{T}/\mathcal{S}$
given by $T\mapsto\mathbb{L}(T)$.
Let us recall the inverse map of the bijection following the proof of [AI, Theorem 2.37]. Denote by $\mathcal{S}_{D}^{\leq 0}:=\cup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D*\operatorname{\mathsf{add}}\nolimits D[1]*\cdots*\operatorname{\mathsf{add}}\nolimits D[l]$ and $\mathcal{S}_{D}^{<0}:=\mathcal{S}_{D}^{\leq 0}[1]$. It is known that $\mathcal{S}_{D}^{<0}$ is covariantly finite in $\mathcal{T}$. Since $\mathcal{S}$ is functorially finite, we may identify $\mathcal{T}/\mathcal{S}$ with $\mathcal{S}^{\perp}$. Let $N\in\mathcal{S}^{\perp}$ be a silting object of $\mathcal{S}^{\perp}$. Consider the following triangle
$S_{N}\to T_{N}\to N\xrightarrow{g}S_{N}[1],$
where $S_{N}[1]\in\mathcal{S}_{D}^{<0}$ and $g$ is a minimal left $\mathcal{S}_{D}^{<0}$-approximation. According to the proof of [AI, Theorem 2.37], $T_{N}\oplus D$ is a silting object of $\mathcal{T}$ such that $\mathbb{L}(T_{N}\oplus D)=\mathbb{L}(N)$.
2.3. Perpendicular subcategories in hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category, that is, an abelian category satisfying $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{2}(-,-)=0$. For any $E\in\mathcal{H}$, define
$E^{\perp_{[0,1]}}=\{X\in\mathcal{H}~|~\operatorname{\mathsf{Hom}}_{\mathcal{H}}(E,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,X)\}.$
Then we have
Lemma 2.6.
$E^{\perp_{[0,1]}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular $E^{\perp_{[0,1]}}$ is closed under images, kernels and cokernels.
Proof.
It is straightforward to check $E^{\perp_{[0,1]}}$ is closed under images, kernels and cokernels in $\mathcal{H}$, and hence it is an abelian subcategory of $\mathcal{H}$. We only need to show $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$ for any $X,Y\in E^{\perp_{[0,1]}}$. We assume a 2-extension $\varepsilon$ in $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)$ which is represented by the following exact sequence in $E^{\perp_{[0,1]}}$
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{u}$${E_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$${\operatorname{\mathsf{im}}\nolimits f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{v}$
Since $E^{\perp_{[0,1]}}$ is an abelian subcategory of $\mathcal{H}$, we obtain two exact sequences in $E^{\perp_{[0,1]}}$
$0\to Y\to E_{1}\xrightarrow{u}\operatorname{\mathsf{im}}\nolimits f\to 0,$ (2.1)
$0\to\operatorname{\mathsf{im}}\nolimits f\xrightarrow{v}E_{2}\to X\to 0.$ (2.2)
We assume that they represent $\varepsilon_{1}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(\operatorname{\mathsf{im}}\nolimits f,Y)$ and $\varepsilon_{2}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(X,\operatorname{\mathsf{im}}\nolimits f)$ respectively. Note that for $M,N\in E^{\perp_{[0,1]}}$, we have $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(M,N)=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(M,N)$. Applying $\operatorname{\mathsf{Hom}}_{E^{\perp_{[0,1]}}}(X,-)$ to (2.1) yields the following exact sequence
$\cdots\to\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(X,E_{1})\xrightarrow{u_{*}}\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(X,\operatorname{\mathsf{im}}\nolimits f)\xrightarrow{\eta}\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)\to\cdots.$
By the definition of the connecting map $\eta$, we have $\varepsilon=\eta(\varepsilon_{2})$. On the other hand, since $\mathcal{H}$ is hereditary, $u_{*}$ is surjective, and hence $\eta=0$. It follows that $\varepsilon=0$. Since $\varepsilon$ is arbitrary, we have $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$. ∎
Denote by ${\mathcal{D}}^{b}(\mathcal{H})$ the bounded derived category of $\mathcal{H}$. As a generalization of $E^{\perp_{[0,1]}}$ for $E\in\mathcal{H}$, we define the following subcategory for any $D\in{\mathcal{D}}^{b}(\mathcal{H})$
$D^{\perp_{\mathbb{Z}}}=\{M\in\mathcal{H}~|~\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D,M[\mathbb{Z}])=0\}.$
The following is a generalization of Lemma 2.6.
Proposition 2.7.
$D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular, $D^{\perp_{\mathbb{Z}}}$ is closed under images, kernels and cokernels.
Proof.
Write $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{r}$, where each $D_{i}$ is indecomposable. Then $D_{i}=X_{i}[n_{i}]$ for some indecomposable object $X_{i}\in\mathcal{H}$ and some $n_{i}\in\mathbb{Z}$. By definition,
$D^{\perp_{\mathbb{Z}}}=\bigcap D_{i}^{\perp_{\mathbb{Z}}}=\bigcap X_{i}^{\perp_{\mathbb{Z}}}=\bigcap X_{i}^{\perp_{[0,1]}}=(\oplus X_{i})^{\perp_{[0,1]}}.$
The assertion holds by Lemma 2.6. ∎
Proposition 2.8.
Let $D\in{\mathcal{D}}^{b}(\mathcal{H})$. Then there exists a triangle equivalence
$\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\xrightarrow{\sim}(\operatorname{\mathsf{thick}}{D})^{\perp}.$
Proof.
Consider the inclusion $D^{\perp_{\mathbb{Z}}}\hookrightarrow\mathcal{H}$. By Proposition 2.7, this inclusion extends to a triangle functor $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to{\mathcal{D}}^{b}(\mathcal{H})$, which maps a complex to itself. Let $X,Y\in D^{\perp_{\mathbb{Z}}}$, by Proposition 2.7 again, we have
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(X,Y)=\operatorname{\mathsf{Hom}}_{D^{\perp_{\mathbb{Z}}}}(X,Y)=\operatorname{\mathsf{Hom}}_{\mathcal{H}}(X,Y)=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,Y),$
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(X,Y[1])=\operatorname{\mathsf{Ext}}_{D^{\perp_{\mathbb{Z}}}}^{1}(X,Y)=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(X,Y)=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,Y[1]).$
It follows that the induced triangle functor $\iota$ is fully faithful. Thus, we may identify ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ as a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$.
Let $\mathcal{S}=\operatorname{\mathsf{thick}}{D}$. Clearly, a complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ is in $\mathcal{S}^{\perp}$. Conversely, given $X\in\mathcal{S}^{\perp}$, since $\mathcal{H}$ is hereditary, $X\cong\oplus_{i}H^{i}X[-i]\in{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. In particular, the image of $\iota$ is $\mathcal{S}^{\perp}$. Hence, $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}{D})^{\perp}$ is a triangle equivalence. ∎
2.4. Rigid objects and presilting complexes over hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category. Recall that $E\in\mathcal{H}$ is called rigid if $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,E)=0$. A presilting (resp. silting) object $T\in{\mathcal{D}}^{b}(\mathcal{H})$ is also called a presilting (resp. silting) complex over $\mathcal{H}$.
The following fundamental result is due to Happel and Ringel [HR].
Lemma 2.9.
Let $E$ and $F$ be indecomposable objects in $\mathcal{H}$ such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(F,E[1])=0$. Then any nonzero homomorphism $f:E\to F$ is a monomorphism or an epimorphism. In particular, the endomorphism ring of an indecomposable rigid object is a division algebra.
Let $\mathcal{X}$ be a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$ and $X\in\mathcal{X}$ an indecomposable object. A path in $\mathcal{X}$ from $X$ to itself is a cycle in $\mathcal{X}$, that is a sequence of nonzero non-isomorphism between indecomposable objects in $\mathcal{X}$ of the form
$X=X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}X_{2}\to\cdots\xrightarrow{f_{r}}X_{r}=X.$
The following is a consequence of Lemma 2.9 (cf. [F, Lemma 4.2] or [HR, Corollary 4.2] ).
Lemma 2.10.
Let $T$ be an object in ${\mathcal{D}}^{b}(\mathcal{H})$ such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,T[1])=0$. Then the subcategory $\operatorname{\mathsf{add}}\nolimits T$ has no cycle.
Lemma 2.10 yields the following consequence.
Lemma 2.11.
Let $T$ be an object in ${\mathcal{D}}^{b}(\mathcal{H})$ such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,T[1])=0$.
  • (a)
    There is a unique decomposition $T=T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$, where each $T_{i}$ is indecomposable such that $T_{i}\in\mathcal{H}[n_{i}]$ for some $n_{1}\leq n_{1}\leq\cdots\leq n_{r}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for $i<j$.
  • (b)
    Assume moreover that $T$ is a presilting object, then
    $\operatorname{\mathsf{thick}}T=\operatorname{\mathsf{thick}}T_{r}*\operatorname{\mathsf{thick}}T_{r-1}*\cdots*\operatorname{\mathsf{thick}}T_{1}.$
Proof.
Since $\mathcal{H}$ is hereditary, we can assume $T=X_{1}[m_{1}]\oplus X_{2}[m_{2}]\oplus\cdots\oplus X_{t}[m_{t}]$ where every $X_{k}\in\mathcal{H}$, $m_{1}<m_{2}<\cdots<m_{t}$. Then every $X_{k}$ is a rigid object in $\mathcal{H}$. By Lemma 2.10, we can assume $X_{k}=X_{k1}\oplus X_{k2}\oplus\cdots\oplus X_{kr_{k}}$, where every $X_{ki}$ is indecomposable and $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(X_{kj},X_{ki})=0$ for $i<j$. Let $n_{ki}=m_{k}$, $T_{ki}=X_{ki}[n_{ki}]=X_{ki}[m_{k}]$ for $1\leq k\leq t$, $1\leq i\leq r_{k}$. Then the decomposation $T=T_{11}\oplus\cdots\oplus T_{1r_{1}}\oplus\cdots\oplus T_{t1}\oplus\cdots\oplus T_{tr_{t}}$ satisfies the assertion of (a).
We now turn to the statement $(b)$. We fix the canonical decomposition of $T$ as in $(a)$. Namely, $T=T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$, where each $T_{i}$ is indecomposable such that $T_{i}\in\mathcal{H}[n_{i}]$ for some $n_{1}\leq n_{1}\leq\cdots\leq n_{r}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for $i<j$. Since $T$ is presilting, we have $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i}[\mathbb{Z}])=0$ when $i<j$. Consequently, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(\operatorname{\mathsf{thick}}T_{j},\operatorname{\mathsf{thick}}T_{i})=0$ when $i<j$. According to [AI, Lemma 2.22 (a)], $\mathcal{U}:=\operatorname{\mathsf{thick}}T_{r}*\operatorname{\mathsf{thick}}T_{r-1}*\cdots*\operatorname{\mathsf{thick}}T_{1}$ is closed under summands and hence it is a thick subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$. Since $T\in\mathcal{U}$ and $\mathcal{U}\subseteq\operatorname{\mathsf{thick}}T$, we conclude that $\operatorname{\mathsf{thick}}T=\mathcal{U}$. ∎
Lemma 2.12.
Let $T$ be a presilting object in ${\mathcal{D}}^{b}(\mathcal{H})$, then $\operatorname{\mathsf{thick}}T$ is functorially finite in ${\mathcal{D}}^{b}(\mathcal{H})$.
Proof.
We fix a decomposition of $T=T_{1}\oplus\cdots\oplus T_{r}$ as in Lemma 2.11 (a). Since $T_{i}$ is indecomposable, it follows that the indecomposable objects of $\operatorname{\mathsf{thick}}T_{i}$ are precisely $T_{i}[j],j\in\mathbb{Z}$. Consequently, $\operatorname{\mathsf{thick}}T_{i}$ is functorially finite for each $1\leq i\leq r$. By Lemma 2.1, we conclude that $\operatorname{\mathsf{thick}}T$ is functorially finite. ∎
Proposition 2.13.
Let $D$ be a presilting object in ${\mathcal{D}}^{b}(\mathcal{H})$, then there is a triangle equivalence
${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\cong{\mathcal{D}}^{b}(\mathcal{H})/\operatorname{\mathsf{thick}}{D}.$
Proof.
By Lemma 2.12, $\operatorname{\mathsf{thick}}D$ is functorially finite in ${\mathcal{D}}^{b}(\mathcal{H})$. It follows that there is a triangle equivalence $\mathbb{L}|_{(\operatorname{\mathsf{thick}}D)^{\perp}}:(\operatorname{\mathsf{thick}}D)^{\perp}\to{\mathcal{D}}^{b}(\mathcal{H})/\operatorname{\mathsf{thick}}D$ by Lemma 2.3. On the other hand, we have a triangle equivalence $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}D)^{\perp}$ by Proposition 2.8. Taking composition of these two functors, the assertion follows immediately. ∎
2.5. Completion
We retain the notation from the previous subsection. The following result from [DF] indicates that any presilitng object in ${\mathcal{D}}^{b}(\mathcal{H})$ can be completed into a silting complex, see also [BY].
Theorem 2.14.
[DF, Theorem 1.1, 1.2] If ${\mathcal{D}}^{b}(\mathcal{H})$ has a silting object, then every presilting object of ${\mathcal{D}}^{b}(\mathcal{H})$ is a direct summand of a silting object.
From now on, we assume that ${\mathcal{D}}^{b}(\mathcal{H})$ has a silting object. Denote by $K_{0}(\mathcal{H})$ the Grothendieck group of ${\mathcal{D}}^{b}(\mathcal{H})$, which is a free abelian group of finite rank. Denote by $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ the rank of $K_{0}(\mathcal{H})$. We have the following consequence.
Corollary 2.15.
Let $D\in{\mathcal{D}}^{b}(\mathcal{H})$ be a presilting object. Then $D$ is silting if and only if $|D|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$.
Proof.
Assume $D$ is basic, and $D\oplus N$ is a basic silting object by Theorem 2.14. By [AI, Theorem 2.27], $|T|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ for any silting object $T\in{\mathcal{D}}^{b}(\mathcal{H})$. Hence $|D|+|N|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$. Then $|D|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})\iff N=0\iff$ $D$ is silting. ∎
A presilting object $D$ is called almost silting if $|D|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})-1$.
Proposition 2.16.
Let $N$ be an almost silting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Then there exists an indecomposable presilting object $D$, such that $D\oplus N$ is silting and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$.
Proof.
Let $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})=n$. Without loss of generality, we assume that $N$ is basic. By Lemma 2.11 (a), we fix a decomposition $N=N_{1}\oplus N_{2}\oplus\cdots\oplus N_{n-1}$ of $N$, where each $N_{i}$ is indecomposable such that $N_{i}\in\mathcal{H}[m_{i}]$ for some $m_{1}\leq m_{2}\leq\cdots\leq m_{n-1}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N_{j},N_{i})=0$ for $i<j$.
Let $T=N\oplus X$ be a basic silting object of ${\mathcal{D}}^{b}(\mathcal{T})$. Define
$p_{T}=\max\{l+1\mid\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{i}[\mathbb{Z}])=0,1\leq i\leq l\}.$
If $p_{T}=n$, then $N\in(\operatorname{\mathsf{thick}}X)^{\perp}$. If $p_{T}<n$, it suffices to show that there is a silting object $M$ with $N$ as a direct summand such that $p_{M}>p_{T}$.
Now assume that $p:=p_{T}<n$. Let
$X\xrightarrow{f}N^{\prime}\to Y\to X[1]$ (2.3)
be the triangle associated with the left mutation of $T$ at $X$. By the definition of $p$, $N^{\prime}\in\operatorname{\mathsf{add}}\nolimits N_{p}\oplus\cdots\oplus N_{n-1}$. It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N^{\prime},N_{i}[\mathbb{Z}])=0$ for any $1\leq i<p$. For any $1\leq i<p$, applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,N_{i}[\mathbb{Z}])$ to the triangle (2.3) yields an exact sequence
$0=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X[1],N_{i}[\mathbb{Z}])\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{i}[\mathbb{Z}])\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N^{\prime},N_{i}[\mathbb{Z}])=0.$
Consequently, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{i}[\mathbb{Z}])=0$ and hence $p_{\mu_{X}^{-}(T)}\geq p$.
If $N_{p}\in\operatorname{\mathsf{add}}\nolimits N^{\prime}$, then $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{p})\neq 0$ and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N_{p},Y)\neq 0$. Since $Y\oplus N_{p}$ is a presilting object, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{p})=0$ by Lemma 2.10. Moreover, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{p}[>0])=0$. Note that $N_{p}\in\mathcal{H}[m_{p}]$ and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N_{p},Y)\neq 0$, it follows that $Y\in\mathcal{H}[m_{p}]$ or $\mathcal{H}[m_{p}+1]$. We conclude that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{p}[<0])=0$, and hence $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{p}[\mathbb{Z}])=0$. Therefore $p_{\mu_{X}^{-}(T)}\geq p+1>p_{T}$ and we may take $M=\mu_{X}^{-}(T)$.
Now assume that $N_{p}\not\in\operatorname{\mathsf{add}}\nolimits N^{\prime}$. Without loss of generality, we assume that $N^{\prime}\in\operatorname{\mathsf{add}}\nolimits N_{k}\oplus\cdots\oplus N_{n-1}$ and $N_{k}\in\operatorname{\mathsf{add}}\nolimits N^{\prime}$ for some $p<k\leq n-1$. It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{p})=0$ and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{k})\neq 0$. By the definition of $p_{T}$, we know that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{p}[-u])\neq 0$ for some $u\geq 1$. Consequently, $X\in\mathcal{H}[m_{p}-u]$ or $\mathcal{H}[m_{p}-u-1]$. On the other hand, by $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X,N_{k})\neq 0$, we have $X\in\mathcal{H}[m_{k}]$ or $\mathcal{H}[m_{k}-1]$. Noticing that $m_{p}\leq\cdots\leq m_{k}$, we conclude that $m_{p}=\cdots=m_{k}$, $u=1$ and $X\in\mathcal{H}[m_{p}-1]$. Consequently, $Y\in\mathcal{H}[m_{p}]$ and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(Y,N_{p})\cong\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(X[1],N_{p})\neq 0$. Now consider the left mutation of $N\oplus Y$ at $Y$. By the above discussion, we conclude that $p_{\mu_{Y}^{-}(N\oplus Y)}>p$ and we may take $M=\mu_{Y}^{-}(N\oplus Y)$. This completes the proof. ∎
The following is a generalization of Proposition 2.16.
Theorem 2.17.
Let $N$ be a presilting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Then there exists a presilting object $D$, such that $D\oplus N$ is silting and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. Moreover, $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$. In particular, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$.
Proof.
Let $N\oplus U$ be a silting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Let $U_{1}$ be an indecomposable direct summand of $U$. By Proposition 2.16, there is a presilting object $D_{1}$ such that $N\oplus(U/U_{1})\oplus D_{1}$ is silting and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{1},N\oplus(U/U_{1})[\mathbb{Z}])=0$. Repeating this operation for other indecomposable direct summands of $U$, we obtain a silting object $N\oplus D_{1}\oplus D_{2}\oplus\cdots\oplus D_{|U|}$, which satisfies $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{i},N[\mathbb{Z}])=0$ for $1\leq i\leq|U|$. Let $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{|U|}$, and then $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$.
Using the same discussion as the proof of Lemma 2.11 (b), we have $\operatorname{\mathsf{thick}}(N\oplus D)=\operatorname{\mathsf{thick}}D*\operatorname{\mathsf{thick}}N$, while $\operatorname{\mathsf{thick}}(N\oplus D)={\mathcal{D}}^{b}(\mathcal{H})$. This means that $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$ and hence $\operatorname{\mathsf{thick}}N=(\operatorname{\mathsf{thick}}D)^{\perp}$. ∎
3. Reduction of endomorphism algebras
In this section, we investigate properties of endomorphism algebras of silting objects over hereditary abelian categories under operations related to idempotents and $\tau$-reduction.
Let $\mathcal{H}$ be a hereditary abelian category. Recall that a rigid object $T\in\mathcal{H}$ is called tilting if $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(T,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(T,X)$ implies $X=0$. A complex $T$ in ${\mathcal{D}}^{b}(\mathcal{H})$ is called a 2-term silting complex over $\mathcal{H}$ ([BZb, Definition 4.1 and Corollary 4.12]) if it satisfies:
  • (S1)
    $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,M[i])=0$ for any $M$ in $\mathcal{H}$ and $i\neq 0$ or $1$.
  • (S2)
    $T$ is a silting object in ${\mathcal{D}}^{b}(\mathcal{H})$.
A $k$-algebra AA is called quasi-tilted if there exists a hereditary abelian category $\mathcal{H}$ and a basic tilting object $T\in\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{\mathcal{H}}(T)$. A $k$-algebra AA is called quasi-silted if there exists a hereditary abelian category $\mathcal{H}$ and a basic 2-term silting complex $T$ over $\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. An artin algebra $\Lambda$ is called shod if for any indecomposable $\Lambda$-module $M$, we have either $\operatorname{\mathsf{pd}}_{\Lambda}M\leq 1$ or $\operatorname{\mathsf{id}}_{\Lambda}M\leq 1$. When $\Lambda$ is shod, a basic property is that $\operatorname{\mathsf{gl.dim}}\Lambda\leq 3$. In particular, a shod algebra $\Lambda$ is called strictly shod if $\operatorname{\mathsf{gl.dim}}\Lambda=3$. The following result from [BZb] reveals the relationship between quasi-silted algebras and shod algebras.
Theorem 3.1.
Let $A$ be a connected finite dimensional algebra over an algebraically closed field. Then $A$ is shod if and only if it is a quasi-silted algebra.
3.1. Idempotent quotients
For a triangulated category $\mathcal{T}$ and a subcategory $\mathcal{X}$, we denote by $\mathcal{X}]$ the ideal of $\mathcal{T}$ consisting of the morphisms in $\mathcal{T}$ which factorize through objects in $\mathcal{X}$.
Lemma 3.2.
Let $T=D\oplus N$ be a basic silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$, and $e_{D}$ be the idempotent of $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ corresponding to $D$. Then $\langle e_{D}\rangle=[\operatorname{\mathsf{add}}\nolimits D](T,T)=[\operatorname{\mathsf{thick}}D](T,T)$.
Proof.
Note that $\langle e_{D}\rangle$ consists of those morphisms in $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ that factor through objects in $\operatorname{\mathsf{add}}\nolimits D$. In other words, $\langle e_{D}\rangle=[\operatorname{\mathsf{add}}\nolimits D](T,T)$. Let $f\in[\operatorname{\mathsf{thick}}D](T,T)$, and assume $f$ factors as $T\xrightarrow{a}X\xrightarrow{b}T$ with $X\in\operatorname{\mathsf{thick}}D$. By Lemma 2.4, we have
$\operatorname{\mathsf{thick}}D=\bigcup\limits_{l\geq 0}\operatorname{\mathsf{add}}\nolimits(D[-l])*\operatorname{\mathsf{add}}\nolimits(D[-l+1])*\cdots*\operatorname{\mathsf{add}}\nolimits(D[l]).$
Assume $l>0$ such that $X\in\operatorname{\mathsf{add}}\nolimits(D[-l])*\operatorname{\mathsf{add}}\nolimits(D[-l+1])*\cdots*\operatorname{\mathsf{add}}\nolimits(D[l])$. Then there exists a triangle $X_{1}\xrightarrow{u}X\to X_{2}\to X_{1}[1]$ with $X_{1}\in\operatorname{\mathsf{add}}\nolimits(D[-l])*\operatorname{\mathsf{add}}\nolimits(D[-l+1])*\cdots*\operatorname{\mathsf{add}}\nolimits D$ and $X_{2}\in\operatorname{\mathsf{add}}\nolimits(D[1])*\operatorname{\mathsf{add}}\nolimits(D[2])*\cdots*\operatorname{\mathsf{add}}\nolimits(D[l])$. Since $T$ is silting, then $\operatorname{\mathsf{Hom}}(T,X_{2})=0$ and hence $a=uc$ for some $c:T\to X_{1}$. There also exists a triangle $X_{3}\to X_{1}\xrightarrow{v}X_{4}\to X_{3}[1]$ with $X_{3}\in\operatorname{\mathsf{add}}\nolimits(D[-l])*\operatorname{\mathsf{add}}\nolimits(D[-l+1])*\cdots*\operatorname{\mathsf{add}}\nolimits(D[-1])$ and $X_{4}\in\operatorname{\mathsf{add}}\nolimits D$. Again since $T$ is silting, then $\operatorname{\mathsf{Hom}}(X_{3},T)=0$ and hence $bu=dv$ for some $d:X_{4}\to T$. Then $f=ba=buc=dvc\in[\operatorname{\mathsf{add}}\nolimits D](T,T)$. ∎
Theorem 3.3.
Let $A$ be an endomorphism algebra of a silting complex over a hereditary abelian category, and $e$ be an idempotent element of $A$.
  • (a)
    Then $A/\langle e\rangle$ is also an endomorphism algebra of a silting complex over a hereditary abelian category.
  • (b)
    Moreover, if $A$ is a quasi-silted algebra, then $A/\langle e\rangle$ is also a quasi-silted algebra.
Proof.
Let $\mathcal{H}$ be a hereditary abelian category, and $T\in{\mathcal{D}}^{b}(\mathcal{H})$ a basic silting object such that $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. Let $T=D\oplus N$ and $e_{D}$ be the idempotent of $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ corresponding to $D$. It suffices to show that $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)/\langle e_{D}\rangle$ is isomorphic to an endomorphism algebra of a basic silting complex over a hereditary abelian category.
Let $\mathcal{S}:=\operatorname{\mathsf{thick}}D$. Since $\mathcal{S}$ is functorially finite, we consider the following triangle in ${\mathcal{D}}^{b}(\mathcal{H})$
$S\xrightarrow{f}N\xrightarrow{g}S_{N}\to S[1],$ (3.4)
where $f$ is a minimal right $\mathcal{S}$-approximation of $N$. It follows that $S_{N}\in\mathcal{S}^{\perp}$ by Lemma 2.2. On the other hand, by Lemma 2.4, it is not hard to see that $S\in\mathcal{S}_{D}^{\leq 0}:=\bigcup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D\ast\cdots\ast\operatorname{\mathsf{add}}\nolimits D[l]$.
Applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,-)$ to the triangle (3.4) yields an exact sequence
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S)\xrightarrow{f_{*}}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,N)\xrightarrow{g_{*}}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S_{N})\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S[1]).$
Since $S\in\mathcal{S}_{D}^{\leq 0}$, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S[1])=0$. By Lemma 3.2, we obtain
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S_{N})\cong\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,N)/\operatorname{\mathsf{im}}\nolimits f_{*}=\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)/\langle e_{D}\rangle.$
On the other hand, applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,S_{N})$ to the triangle (3.4), we obtain the following exact sequence
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S[1],S_{N})\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{N},S_{N})\xrightarrow{g^{\ast}}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S_{N})\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S,S_{N}),$
where $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S[1],S_{N})=0=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S,S_{N})$ since $S_{N}\in\mathcal{S}^{\perp}$. Consequently, we obtain isomorphisms of vector spaces
$\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{N},S_{N})\xrightarrow{g^{\ast}}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,S_{N})\xleftarrow{\bar{g}_{*}}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,N)/\operatorname{\mathsf{im}}\nolimits f_{*}.$
It is routine to check that $\bar{g}_{*}^{-1}\circ g^{\ast}$ is an isomorphism of algebras.
Let $\mathbb{L}:{\mathcal{D}}^{b}(\mathcal{H})\to{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$ be the localization functor. According to Proposition 2.13, we have ${\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}\cong{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Clearly, $\mathbb{L}(T)=\mathbb{L}(S_{N})$. By Theorem 2.5, $\mathbb{L}(S_{N})$ is a basic silting object of ${\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$. Since $S_{N}\in\mathcal{S}^{\perp}$, we conclude that $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(\mathbb{L}(N))\cong\operatorname{\mathsf{End}}_{\mathcal{S}^{\perp}}(S_{N})\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{N})$ by Lemma 2.3, and hence $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(\mathbb{L}(N))\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)/\langle e_{D}\rangle$. This finishes the proof of $(a)$.
Now assume $T$ is 2-term (i.e. , $T$ satisfies (S1)) in ${\mathcal{D}}^{b}(\mathcal{H})$. By Proposition 2.8, we may identify ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ with $\mathcal{S}^{\perp}$ via the embedding $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\hookrightarrow{\mathcal{D}}^{b}(\mathcal{H})$. It suffices to verify that $S_{N}$ is 2-term in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$.
Let $M\in D^{\perp_{\mathbb{Z}}}$ and $i\neq 0,1$. Applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,M[i])$ to the triangle (3.4), we obtain an exact sequence
$0\overset{M\in D^{\perp_{\mathbb{Z}}}}{=}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S,M[i-1])$ $\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{N},M[i])$
$\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N,M[i])\overset{T~\text{is 2-term}}{=}0.$
It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(S_{N},M[i])=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{N},M[i])=0$ for $i\neq 0,1$. Hence $S_{N}$ is a 2-term silting complex over $D^{\perp_{\mathbb{Z}}}$. ∎
Remark 3.4.
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $e$ be an idempotent element in $A$. By Theorem 3.1 and Theorem 3.3 (b), if $A$ is shod, then $A/\langle e\rangle$ is also shod.
3.2. Idempotent subalgebras
Theorem 3.5.
Let $A$ be an endomorphism algebra of a basic silting complex over a hereditary category, and $e$ be an idempotent element in $A$.
  • (a)
    The subalgebra $eAe$ is also an endomorphism algebra of a basic silting complex over a hereditary category.
  • (b)
    If $A$ is a quasi-silted algebra, then $eAe$ is also a quasi-silted algebra.
  • (c)
    If $A$ is a quasi-tilted algebra, then $eAe$ is also a quasi-tilted algebra.
Proof.
(a) Assume $A=\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$, where $\mathcal{H}$ is a hereditary abelian category, and $T$ is a basic silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Assume $T=U\oplus N$ and $e=e_{N}$ is the idempotent corresponding to $N$. We first note that $e_{N}(\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T))e_{N}\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)$. By Theorem 2.17, there exists a presilting complex $D$, such that $D\oplus N$ is a basic silting object and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. In addition, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$. By Proposition 2.7, $D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian category. By Proposition 2.8, we identify $(\operatorname{\mathsf{thick}}D)^{\perp}$ with ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Then $N$ is a basic silting complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$, and $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{D^{\perp_{\mathbb{Z}}}})}(N)$.
(b) Assume $T$ is 2-term in ${\mathcal{D}}^{b}(\mathcal{H})$. Since $N$ is a direct summand of $T$, then $N$ is 2-term in ${\mathcal{D}}^{b}(\mathcal{H})$, and hence $N$ is 2-term in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$.
(c) Assume $T$ is tilting in $\mathcal{H}$. Then $N\in D^{\perp_{\mathbb{Z}}}$. By [BZb, Proposition 4.8], $N$ is tilting in $D^{\perp_{\mathbb{Z}}}$. ∎
Remark 3.6.
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $e$ be an idempotent element in $A$. By Theorem 3.1 and Theorem 3.5 (b), if $A$ is shod, then $eAe$ is also shod.
Remark 3.7.
The statements in parts (b) and (c) of Theorem 3.5 was established in [AC]. Our method is different with the one in [AC].
$3.$ $\tau$-tilting reduction
Let $A$ be a finite-dimensional $k$-algebra and $\operatorname{\mathsf{mod}}\nolimits A$ the category of finite dimensional right $A$-modules. Recall that $M\in\operatorname{\mathsf{mod}}\nolimits A$ is $\tau$-rigid if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$, where $\tau$ is the Auslander–Reiten translation functor. A $\tau$-rigid module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is $\tau$-tilting if $|M|=|A|$. A $\tau$-rigid module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called a support $\tau$-tilting $A$-module, if there is an idempotent $e\in A$ such that $M$ is a $\tau$-tilting $A/\langle e\rangle$-module. Denote by $s\tau\text{-}\operatorname{\mathsf{tilt}}A$ the set of isomorphism classes of basic support $\tau$-tilting $A$-modules.
Let $Z\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid $A$-module, then $^{\perp}\tau Z:=\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,\tau Z)=0\}$ is a torsion class of $\operatorname{\mathsf{mod}}\nolimits A$. The Bongartz completion $U_{Z}:=P(^{\perp}(\tau Z))$ of $Z$ is the maximal basic direct sum of $\operatorname{\mathsf{Ext}}$-projective objects in $^{\perp}(\tau Z)$, which is a $\tau$-tilting module. Let $B=\operatorname{\mathsf{End}}_{A}(U_{Z})$ and $C=B/\langle e_{Z}\rangle$, where $e_{Z}$ is the idempotent of $B$ corresponding to the projective $B$-module $\operatorname{\mathsf{Hom}}_{A}(U_{Z},Z)$. The algebra $C$ is called the $\tau$-tilting reduction of $A$ with respect to $Z$, see [J].
Before stating the main result of this subsection. Let us recall a result from [J]. Let $\mathcal{T}$ be a $k$-linear $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulate category with a silting object $T$. Let $A=\operatorname{\mathsf{End}}_{\mathcal{T}}(T)$, and denote
$\overline{(-)}=\operatorname{\mathsf{Hom}}_{\mathcal{T}}(T,-):\mathcal{T}\to\operatorname{\mathsf{mod}}\nolimits A.$
Let $D$ be a presilting object in $\mathcal{T}$ such that $D\in\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits T[1]$. Let $f:D^{\prime}\to T[1]$ be a minimal right $(\operatorname{\mathsf{add}}\nolimits D)$-approximation of $T[1]$. By embedding $f$ into a distinguished triangle, we obtain
$T\to X_{D}\to D^{\prime}\xrightarrow{f}T[1].$ (3.5)
Then $T_{D}=X_{D}\oplus D$ lies in 2T-$\operatorname{\mathsf{silt}}\mathcal{T}:=\{T\in\operatorname{\mathsf{silt}}\mathcal{T}~|~T\in\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits(T[1])\}$ (see [J, Proposition 4.9]). Jasso [J] has established the following connection between silting reduction and $\tau$-tilting reduction.
Theorem 3.8.
[J, Theorem 4.12 (a)] Let $D$ be a presilting object in $\mathcal{T}$ contained in $\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits T[1]$. Then we have $\operatorname{\mathsf{End}}_{A}(U_{\overline{D}})/\langle e_{\overline{D}}\rangle\cong\operatorname{\mathsf{End}}_{\mathcal{T}/\mathcal{S}}(T_{D})$, where $\mathcal{S}=\operatorname{\mathsf{thick}}D$ and $e_{\overline{D}}$ is the idempotent of $\operatorname{\mathsf{End}}_{A}(U_{\overline{D}})$ corresponding to the projective $\operatorname{\mathsf{End}}_{A}(U_{\overline{D}})$-module $\operatorname{\mathsf{Hom}}_{A}(U_{\overline{D}},\overline{D})$.
Now we are in the position to state the main result of this subsection.
Theorem 3.9.
Let $A$ be an endomorphism algebra of a basic silting complex over a hereditary abelian category and $Z$ be a $\tau$-rigid $A$-module.
  • (a)
    The $\tau$-tilting reduction of $A$ with respect to $Z$ is an endomorphism algebra of a basic silting complex over a hereditary abelian category.
  • (b)
    If $A$ is a quasi-silted algebra, then the $\tau$-tilting reduction of $A$ with respect to $Z$ is also a quasi-silted algebra.
Proof.
[…]
For the statement $(b)$. Let assume moreover $T$ is a 2-term silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Consider the following triangle in ${\mathcal{D}}^{b}(\mathcal{H})$
$S\xrightarrow{f}T\to S_{T}\to S[1],$
where $f$ is a minimal right $\mathcal{S}$-approximation of $T$. We have $S_{T}\in\mathcal{S}^{\perp}$. We identify $\mathcal{S}^{\perp}$ with ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ by Proposition 2.8. It follows that
$\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(T_{D})\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(T)\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(S_{T}).$
It suffices to show that $S_{T}$ is a 2-term silting complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$.
Note that $S_{T}\cong T$ is silting in ${\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$ by Theorem 2.5. Then $S_{T}$ is silting in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Let $M\in D^{\perp_{\mathbb{Z}}}$ and $i\neq 0,1$. Applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,M[i])$ to the above triangle, we obtain an exact sequence
$0\overset{M\in D^{\perp_{\mathbb{Z}}}}{=}\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S,M[i-1])$ $\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{T},M[i])$
$\to\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,M[i])\overset{T~\text{is 2-term}}{=}0.$
It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(S_{T},M[i])=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{T},M[i])=0$ for $i\neq 0,1$. Hence $S_{N}$ is a 2-term silting complex over $D^{\perp_{\mathbb{Z}}}$. This completes the proof. ∎
Remark 3.10.
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $Z$ be a $\tau$-rigid $A$-module. By Theorem 3.1 and Theorem 3.9 (b), if $A$ is shod, then the $\tau$-tilting reduction $\operatorname{\mathsf{End}}_{A}(U_{Z})/\langle e_{Z}\rangle$ is also a shod algebra.
Remark 3.11.
For a shod algebra $A$, and a $\tau$-rigid $A$-module $Z$, the endomorphism algebra $\operatorname{\mathsf{End}}_{A}(U_{Z})$ may not be shod, see Example 5.4.
4. Idempotent quotients for several classes of algebras
This section is devoted to proving the following result.
Theorem 4.1.
Let $A$ be a finite-dimensional $k$-algebra and $e$ an idempotent element in $A$.
  1. (1)
    If $A$ is a laura algebra, then so is $A/\langle e\rangle$.
  2. (2)
    If $A$ is a right (or left) glued algebra, then so is $A/\langle e\rangle$.
  3. (3)
    If $A$ is a weakly shod algebra, then so is $A/\langle e\rangle$.
  4. (4)
    If $A$ is a shod algebra, then so is $A/\langle e\rangle$.
Remark 4.2.
Theorem 4.1 (4) also follows from Theorem 3.3 (b).
For an artin algebra $A$, denote by $\operatorname{\mathsf{ind}}A$ the set of isomorphism classes of indecomposable right $A$-modules. Given indecomposable $A$-modules $M$ and $N$, a path from $M$ to $N$ is a sequence
$M=X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}X_{2}\to\cdots\xrightarrow{f_{s}}X_{s}=N$
where every $X_{i}\in\operatorname{\mathsf{ind}}A$ and every $f_{i}$ is a nonzero non-isomorphic morphism. The number $s$ is called the length of this path. Following [HRS], we have two full subcategories
$\mathcal{L}_{A}=\{X\in\operatorname{\mathsf{ind}}A~|~\text{for any path from }~W~\text{to}~X,~\text{we have}~\operatorname{\mathsf{pd}}_{A}W\leq 1\},$
$\mathcal{R}_{A}=\{Y\in\operatorname{\mathsf{ind}}A~|~\text{for any path from }~Y~\text{to}~W,~\text{we have}~\operatorname{\mathsf{id}}_{A}W\leq 1\}.$
We recall that
  • AA is called laura if $\mathcal{L}_{A}\cup\mathcal{R}_{A}$ is cofinite in $\operatorname{\mathsf{ind}}A$, that is, there are only finitely many modules in $\operatorname{\mathsf{ind}}A-\mathcal{L}_{A}\cup\mathcal{R}_{A}$;
  • AA is called right glued (resp. left glued) if $\mathcal{L}_{A}$ (resp. $\mathcal{R}_{A}$) is cofinite in $\operatorname{\mathsf{ind}}A$ (see [AC, Lemma 3.1]);
  • AA is called weakly shod if the length of any path from an indcomposable injective module to an indcomposable projective module is bounded.
Proposition 4.3.
[AC, Proposition 1.3] An artin algebra is weakly shod if and only if there exists an $l\geq 0$ such that any path from an indecomposable module not lying in $\mathcal{L}_{A}$ to an indecomposable module not lying in $\mathcal{R}_{A}$ has length at most $l$.
Henceforth, we assume that $A$ is a finite-dimensional $k$-algebra.
Lemma 4.4.
Let $e$ be an idempotent in $A$, for a right $A$-module $M$, we have a short exact sequence $0\to MeA\to M\to M/MeA\to 0$, where $MeA\in\operatorname{\mathsf{Fac}}eA$ and $M/MeA\in(eA)^{\perp}$.
Proof.
Consider a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$, $f:Q\to M$. Then we have a short exact sequence $0\to\operatorname{\mathsf{im}}\nolimits f\to M\xrightarrow{g}M/\operatorname{\mathsf{im}}\nolimits f\to 0$. Then $\operatorname{\mathsf{im}}\nolimits f\in\operatorname{\mathsf{Fac}}eA$. Set a morphism $h:eA\to M/\operatorname{\mathsf{im}}\nolimits f$. Since $eA$ is projective, $h$ is factorized by $g$, say $h=gu$ for some $u:eA\to M$. Then $u$ is factorized by $f$, since $f$ is a right $\operatorname{\mathsf{add}}\nolimits eA$-approximation. Then $u=fv$ for some $v:eA\to Q$. Hence $h=gu=gfv=0$. This means $M/\operatorname{\mathsf{im}}\nolimits f\in(eA)^{\perp}$.
Next we show $\operatorname{\mathsf{im}}\nolimits f=MeA$. Note that $\operatorname{\mathsf{im}}\nolimits f$ is independent of the choice of the right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation $f$. Hence firstly we construct a certain right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$. Assume $\{x_{1}e,\cdots,x_{d}e\}$ is a basis of the $k$-vector space $Me$. For each $1\leq i\leq d$, denote by $(x_{i}e)\cdot$ the left multiplication map $eA\to M$, $ea\mapsto x_{i}ea$, which is a right $A$-morphism. Then we have a morphism $\phi:(eA)^{d}\to M$ by the universal property of direct sum. By the isomorphism $Me\cong\operatorname{\mathsf{Hom}}_{A}(eA,M)$, $\{(x_{1}e)\cdot,\cdots,(x_{d}e)\cdot\}$ is a basis of $\operatorname{\mathsf{Hom}}_{A}(eA,M)$. Hence $\phi$ is a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$. In particular, $\operatorname{\mathsf{im}}\nolimits f=\operatorname{\mathsf{im}}\nolimits\phi$. By the definition of $\phi$, for any $\{ea_{i}\}_{i}\in(eA)^{d}$, $\phi(\{ea_{i}\}_{i})=x_{1}ea_{1}+\cdots+x_{d}ea_{d}$. This means $\operatorname{\mathsf{im}}\nolimits\phi\subseteq MeA$. On the other hand, $MeA$ is generated by elements in form of xeaxea with $x\in M$, $a\in A$. Set $xe=\sum\limits_{i=1}^{d}\lambda_{i}x_{i}e$, $\lambda_{i}\in k$. Then $xea=\sum\limits_{i=1}^{d}x_{i}e(\lambda_{i}a)\in\operatorname{\mathsf{im}}\nolimits\phi$. Hence $MeA=\operatorname{\mathsf{im}}\nolimits\phi=\operatorname{\mathsf{im}}\nolimits f$. This finishes the proof. ∎
Remark 4.5.
The projective $A$-module eAeA induces a torsion pair $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$ of $\operatorname{\mathsf{mod}}\nolimits A$. The sequence in Lemma 4.4 is the canonical sequence for $M$ with respect to $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$.
Lemma 4.6.
For $M\in(eA)^{\perp}$, M is $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$, if and only if there is a projective $A$-module $P$ such that $M\cong P/PeA$.
Proof.
Note that $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$ is a torsion pair in $\operatorname{\mathsf{mod}}\nolimits A$. By Lemma 4.4 and [ASS, Proposition VI.1.11 (b)], the assertion holds. ∎
Lemma 4.7.
For the standard duality $\operatorname{\mathsf{D}}\nolimits=\operatorname{\mathsf{Hom}}_{k}(-,k)$, we have the following properties:
  1. (1)
    For an $A$-module $X$, we have $\operatorname{\mathsf{id}}_{A}X=\operatorname{\mathsf{pd}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$ and $\operatorname{\mathsf{pd}}_{A}X=\operatorname{\mathsf{id}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$. In particular, $A$ is shod if and only if $A^{op}$ is shod.
  2. (2)
    $(A/\langle e\rangle)^{op}=A^{op}/\langle e\rangle$.
  3. (3)
    For a $A$-module $X$, $X\in(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)\in(Ae)^{\perp}$. In particular, the restriction of $\operatorname{\mathsf{D}}\nolimits$ to $(eA)^{\perp}$ is a duality, $\operatorname{\mathsf{D}}\nolimits:(eA)^{\perp}\to(Ae)^{\perp}$.
  4. (4)
    For a $A$-module $X\in(eA)^{\perp}$, it is $\operatorname{\mathsf{Ext}}$-projective (resp. $\operatorname{\mathsf{Ext}}$-injective) in $(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)$ is $\operatorname{\mathsf{Ext}}$-injective (resp. $\operatorname{\mathsf{Ext}}$-projective) in $(Ae)^{\perp}$.
Proof.
(1) and (2) are clear.
For (3), take an $A$-module $X$. Then
$X\in(eA)^{\perp}$ $\iff\operatorname{\mathsf{Hom}}_{A}(eA,X)=0$
$\iff\operatorname{\mathsf{Hom}}_{A^{op}}(\operatorname{\mathsf{D}}\nolimits(X),\operatorname{\mathsf{D}}\nolimits(eA))=0$
$\iff\operatorname{\mathsf{Hom}}_{A^{op}}(Ae,\operatorname{\mathsf{D}}\nolimits(X))=0$
$\iff\operatorname{\mathsf{D}}\nolimits(X)\in(Ae)^{\perp}.$
Statement (4) follows from (3) by the duality $\operatorname{\mathsf{Ext}}_{A}^{1}(X,Y)\cong\operatorname{\mathsf{Ext}}_{A^{\text{op}}}^{1}(\operatorname{\mathsf{D}}\nolimits(Y),\operatorname{\mathsf{D}}\nolimits(X)$, which holds for any $X,Y\in\operatorname{\mathsf{mod}}\nolimits A$. ∎
The following is well-known.
Lemma 4.8.
Let $A$ be a finite-dimensional $k$-algebra, II an ideal of $A$ and $B=A/I$, then the following properties hold.
  1. (1)
    $\operatorname{\mathsf{mod}}\nolimits B$ can be viewed as a subcategory of $\operatorname{\mathsf{mod}}\nolimits A$. In particular, a $B$-module is indecomposable if and only if it is an indecomposable $A$-module.
  2. (2)
    $\operatorname{\mathsf{mod}}\nolimits B$ is closed under factor modules, submodules, kernels and cokernels in $\operatorname{\mathsf{mod}}\nolimits A$.
  3. (3)
    Given a morphism $f:X\to Y$ in $\operatorname{\mathsf{mod}}\nolimits B$, then $f$ is monomorphic (resp. epimorphic) in $\operatorname{\mathsf{mod}}\nolimits B$ if and only if it is monomorphic (resp. epimorphic) in $\operatorname{\mathsf{mod}}\nolimits A$.
  4. (4)
    A sequence $0\to X\to Y\to Z\to 0$ in $\operatorname{\mathsf{mod}}\nolimits B$ is exact if and only if it is exact when viewed as a sequence in $\operatorname{\mathsf{mod}}\nolimits A$.
Proposition 4.9.
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$, if $\operatorname{\mathsf{pd}}_{A}M\leq 1$, then $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\leq 1$.
Proof.
Assume $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, and we prove $\operatorname{\mathsf{pd}}_{A}M\geq 2$. Note that for an $A$-module in $(eA)^{\perp}=\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, it is $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$ if and only if it is a projective $A/\langle e\rangle$-module. Let $0\to L\to N\to M\to 0$ be a short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, where $N$ is a projective $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$-module. By Lemma 4.6, $N\cong P/PeA$ for some projective $A$-module $P$. By Lemma 4.8 (4), the above short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$ is also a short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A$. Considering the pull-back of morphisms $L\to N$ and $P\to N$, we have the following commutative diagram in $\operatorname{\mathsf{mod}}\nolimits A$:
[…]
Since $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, $L$ is not $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Note that $PeA\in\operatorname{\mathsf{Fac}}eA$ and $L\in(eA)^{\perp}$. The first column in the above diagram is the canonical sequence of X. Hence, by Lemma 4.4, $L\cong X/XeA$. By Lemma 4.6 again, $X$ is non-projective in $\operatorname{\mathsf{mod}}\nolimits A$ since $L$ is not $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Hence $\operatorname{\mathsf{pd}}_{A}M\geq 2$, we obtain a contradiction. ∎
Proposition 4.10.
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$, if $\operatorname{\mathsf{id}}_{A}M\leq 1$, then $\operatorname{\mathsf{id}}_{A/\langle e\rangle}M\leq 1$.
Proof.
By Lemma 4.7 (1), $\operatorname{\mathsf{pd}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(M)=\operatorname{\mathsf{id}}_{A}M\leq 1$. By Lemma 4.7 (3), $\operatorname{\mathsf{D}}\nolimits(M)\in(Ae)^{\perp}$. By Lemma 4.7 (1), (2) and Proposition 4.9, $\operatorname{\mathsf{id}}_{A/\langle e\rangle}M=\operatorname{\mathsf{pd}}_{A^{op}/\langle e\rangle}\operatorname{\mathsf{D}}\nolimits(M)\leq 1$. ∎
Proposition 4.11.
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$,
  1. (1)
    if $M\in\mathcal{L}_{A}$, then $M\in\mathcal{L}_{A/\langle e\rangle}$;
  2. (2)
    if $M\in\mathcal{R}_{A}$, then $M\in\mathcal{R}_{A/\langle e\rangle}$.
Proof.
We only prove (1); the proof of (2) is similar. Assume there is a path from an indecomposable module $X$ to $M$ in $\operatorname{\mathsf{ind}}A/\langle e\rangle$. Then this is also a path from $X$ to $M$ in $\operatorname{\mathsf{ind}}A$. Since $M\in\mathcal{L}_{A}$, then $\operatorname{\mathsf{pd}}_{A}X\leq 1$. By Proposition 4.9, we have $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}X\leq 1$. This means $M\in\mathcal{L}_{A/\langle e\rangle}$. ∎
With these preparations in hand, we now turn to the proof of Theorem 4.1.
proof of Theorem 4.1.
(1) By Proposition 4.11, for an indecomposable $A/\langle e\rangle$-module $M$, if $M\notin\mathcal{L}_{A/\langle e\rangle}\cup\mathcal{R}_{A/\langle e\rangle}$, then $M\notin\mathcal{L}_{A}\cup\mathcal{R}_{A}$. Hence if $\mathcal{L}_{A}\cup\mathcal{R}_{A}$ is cofinite, so is $\mathcal{L}_{A/\langle e\rangle}\cup\mathcal{R}_{A/\langle e\rangle}$.
The proof of (2) is similar to (1).
(3) By Proposition 4.3, assume $l\geq 0$ is an integer satisfying that any path from an indecomposable module not lying in $\mathcal{L}_{A}$ to an indecomposable module not lying in $\mathcal{R}_{A}$ has length at most $l$. We consider a path $X\to\cdots\to Y$ in $\operatorname{\mathsf{ind}}A/\langle e\rangle$, where $X$, $Y$ are indecomposable and $X\notin\mathcal{L}_{A/\langle e\rangle}$, $Y\notin\mathcal{R}_{A/\langle e\rangle}$. By Proposition 4.11, we have $X\notin\mathcal{L}_{A}$, $Y\notin\mathcal{R}_{A}$. Since this path is also a path in $\operatorname{\mathsf{ind}}A$, its length $\leq l$. By Proposition 4.3 again, we know that $A/\langle e\rangle$ is also weakly shod.
(4) Given an indecomposable $A$-module $M$, we have $\operatorname{\mathsf{pd}}_{A}M\leq 1$ or $\operatorname{\mathsf{id}}_{A}M\leq 1$ since $A$ is shod. According to Proposition 4.9 and Proposition 4.10, this assertion holds immediately. ∎
Remark 4.12.
The statement in Theorem 4.1 does not hold for tilted algebras and quasi-tilted algebras, while, in contrast, the result for special idempotents in [Z] remains valid for these two classes. Indeed, there is an algebra in the example from [BMR, Section 2.5] which is a tilted algebra, becoming strictly shod instead after taking a quotient by an idempotent.
5. Examples
We consider a class of algebras. Let $n\geq 2$, $2\leq k\leq n$. Denote by $A(n,k)$ the bound quiver algebra given by:
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{2}}$${k\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{k-1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{k}}$${n-1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{n-2}}$${n\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{~\alpha_{n-1}}$
with
$\alpha_{2}\alpha_{1}=0,~\alpha_{3}\alpha_{2}=0,~\cdots,~\alpha_{k-1}\alpha_{k-2}=0.$
Here, we adopt the convention for the composition of arrows as established in [ASS]. Clearly, $A(n,2)$ is a hereditary algebra of type $\mathbb{A}_{n}$, and $A(n,n)$ is the quotient of $A(n,2)$ modulo the square of its Jacobson radical.
Fix $n\geq 3$ and $2\leq k<n$. For each vertex $i$ of the quiver, we denote the corresponding simple module by $S(i)$, its projective cover by $P(i)$, and its injective envelop by $I(i)$. The Auslander-Reiten quiver of $A(n,k)$ has the following shape
[…]
$S(1)=P(1)$ is projective. So $\operatorname{\mathsf{pd}}_{A(n,k)}S(1)=0$. For $1<i\leq k$, we have a short exact sequence $0\to S(i-1)\to P(i)\to S(i)\to 0$. By induction, $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=i-1$ for $1\leq i\leq k$. For $i>k$, we have a short exact sequence $0\to P(i-1)\to P(i)\to S(i)\to 0$. So $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=1$ for $k<i\leq n$. Hence $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. (By the same argument, one can check $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$.)
We take $T=P(1)\oplus P(2)\oplus\cdots\oplus P(k-1)\oplus P(n)\oplus I(k+1)\oplus\cdots\oplus I(n)$. Clearly, $|T|=n=|A(n,k)|$ and it is easy to see $\operatorname{\mathsf{pd}}_{A(n,k)}T\leq 1$. A straightforward computation shows that
$\tau I(k+1)={\begin{smallmatrix}n-1\\ \vdots\\ k\end{smallmatrix}},~\tau I(k+2)={\begin{smallmatrix}n-1\\ \vdots\\ k+1\end{smallmatrix}},\cdots,~\tau I(n)={\begin{smallmatrix}n-1\end{smallmatrix}}.$
Since there is no path in this Auslander-Reiten quiver from an injective module to one of these modules, we have
$\operatorname{\mathsf{Hom}}_{A(n,k)}(P(n)\oplus I(k+1)\oplus\cdots\oplus I(n),\tau T)=0.$
On the other hand, comparing the dimension vectors of $P(1),\cdots,P(k-1)$ and $\tau I(k+1),\cdots,\tau I(n)$, it is easy to see
$\operatorname{\mathsf{Hom}}_{A(n,k)}(P(1)\oplus P(2)\oplus\cdots\oplus P(k-1),\tau T)=0.$
Hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T,\tau T)=0$, and then $T$ is a tilting $A(n,k)$-module.
Next we consider the endomorphism algebra $B=\operatorname{\mathsf{End}}_{A(n,k)}T$. Write $T=T_{1}\oplus\cdots\oplus T_{n}$ where $T_{i}=P(i)$ for $1\leq i\leq k-1$, $T_{k}=P(n)$, $T_{i}=I(i)$ for $k+1\leq i\leq n$. Observe the above Auslander-Reiten quiver and the positions of these $T_{i}$. We have the following facts:
  • For $1\leq i<j\leq n$, there is no path from $T_{j}$ to $T_{i}$ and hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{j},T_{i})=0$.
  • For $1\leq i\leq n-1$, there is only 1 path from $T_{i}$ to $T_{i+1}$ which is sectional. Composing the irreducible morphisms in this path, we obtain a nonzero morphism $f_{i}:T_{i}\to T_{i+1}$. Hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{i+1})$ is a 1-dimensional $k$-vector space spanned by $f_{i}$.
  • For $1\leq i\leq k-1$, $i+2\leq j\leq n$, by comparing the dimension vectors of $T_{i}$ and $T_{j}$, we have $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{j})=0$.
  • For $k\leq i\leq n-2$, $i+2\leq j\leq n$, notice that $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{j})$ is a 1-dimensional $k$-vector space spanned by $f_{ij}=f_{j-1}f_{j-2}\cdots f_{i}$.
Denote by $e_{i}$ the primitive idempotent of $B$ corresponding to $T_{i}$. With the above observations, we have
$e_{i}(\operatorname{\mathsf{rad}}B)e_{j}=0,~\text{ for }j>i,$
$e_{i+1}(\operatorname{\mathsf{rad}}B)e_{i}/e_{i+1}(\operatorname{\mathsf{rad}}B)^{2}e_{i}\cong kf_{i},~\text{ for }1\leq i\leq n-1,$
$e_{j}(\operatorname{\mathsf{rad}}B)e_{i}/e_{j}(\operatorname{\mathsf{rad}}B)^{2}e_{i}=0,~\text{ for }1\leq i<j\leq n,j-i\geq 2.$
In addition, $f_{2}f_{1}=0$, $f_{3}f_{2}=0$, $\cdots$, $f_{k}f_{k-1}=0$.
Hence $B$ is given by
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{2}}$${k\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{k-1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{k}}$${n-1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{n-2}}$${n\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{~\alpha_{n-1}}$
with
$\alpha_{2}\alpha_{1}=0,~\alpha_{3}\alpha_{2}=0,~\cdots,~\alpha_{k}\alpha_{k-1}=0.$
This implies $B\cong A(n,k+1)$. We have thus shown the following result.
Proposition 5.1.
For $n\geq 2$ and $2\leq k\leq n$, we have $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. If $k<n$, there is a tilting $A(n,k)$-module $T$ such that $\operatorname{\mathsf{End}}_{A(n,k)}T\cong A(n,k+1)$. In particular, iteratively taking the endomorphism algebras of tilting modules, starting from a hereditary algebra of type $\mathbb{A}_{n}$, yields an algebra of global dimension n1n-1 after n2n-2 steps.
Proof.
It suffices to note that $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$ and $A(n,2)$ is a hereditary algebra of type $\mathbb{A}_{n}$. ∎
Let $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}X_{2}\to\cdots\xrightarrow{f_{s}}X_{s}$ be a path of irreducible morphisms in $\operatorname{\mathsf{ind}}A$. If $\tau X_{i+1}=X_{i-1}$ for some $1\leq i\leq s-1$, then $X_{i}$ is called a hook in this path (cf. [CL]).
We also have the following.
Proposition 5.2.
For $n\geq 2$ and $2\leq k\leq n$, we have
  1. (1)
    $A(n,2)$ is a hereditary algebra.
  2. (2)
    $A(n,3)$ is a quasi-tilted algebra.
  3. (3)
    $A(n,4)$ is a strictly shod algebra.
  4. (4)
    $A(n,k)$ is not shod for $k\geq 5$.
Proof.
(1) is clear. For (4), we note that $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1\geq 4$ for $k\geq 5$. Hence $A(n,k)$ is not shod since the global dimension of a shod algebra is not larger than 3.
Let $k=3$. Observe that, in the Auslander-Reiten quiver of $A(n,3)$, there is a sectional path from $I(1)$ to $P(n)$, and this is the only path from an injective $A(n,3)$-module to a projective $A(n,3)$-module. By [HRS, Theorem II.1.14], $A(n,3)$ is a quasi-tilted algebra.
Let $k=4$. There is a path from $I(1)$ to $P(n)$ and a path from $I(2)$ to $P(n)$. These are the only paths from an injective $A(n,4)$-module to a projective $A(n,4)$-module. The first has one hook and the second is sectional. Hence $A(n,4)$ is a strictly shod algebra by [CL, Theorem 2.1 and Proposition 2.4]. ∎
Example 5.3.
Let $A$ be the $k$-algebra given by:
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${4\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$
with $\beta\alpha=0$. By Proposition 5.2 (2), $A\cong A(4,3)$ is quasi-tilted. Its Auslander-Reiten quiver is in the following shape
[…]
We take $T=P(1)\oplus P(2)\oplus P(4)\oplus I(4)$, which is the Bongartz completion of $I(4)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}T$ is given by:
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${4\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$
with $\beta\alpha=0$, $\gamma\beta=0$. By Proposition 5.2 (3), $B\cong A(4,4)$ is strictly shod. in particular, it is not quasi-tilted again.
Example 5.4.
Let $A$ be the $k$-algebra given by:
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${4\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$${5\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta}$
with $\beta\alpha=0$, $\gamma\beta=0$. By Proposition 5.2 (3), $A\cong A(5,4)$ is strictly shod. Its Auslander-Reiten quiver is
[…]
We take $T=P(1)\oplus P(2)\oplus P(3)\oplus P(5)\oplus I(5)$, which is the Bongartz completion of $I(5)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}T$ is given by:
${1}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${3\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${4\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$${5\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta}$
with $\beta\alpha=0$, $\gamma\beta=0$, $\delta\gamma=0$. It is easy to check $\operatorname{\mathsf{pd}}_{B}S(3)=\operatorname{\mathsf{id}}_{B}S(3)=2$, and hence $B=\operatorname{\mathsf{End}}_{A}T$ is not shod.